OpenStudy (anonymous):

Help please... show that A,B,C are collinear OA=(4,14,-5) OB=(1,2,7) OC= (2,6,3)

OpenStudy (mathmate):

If A,B,C are collinear, unit vectors of AB and BC should be equal (identical), or AB=k*BC where k \(\in\)R. AB=<1-4,2-14,7-(-5))=<-3,-12,12> BC=<2-1,6-2,3-7)=<1,4,-4> I'll leave you to do the second part AB=k*BC.

OpenStudy (phi):

collinear means "lie on the same line" I would show all 3 points lie on the line P= (1,2,7)+ t*(1,4,-4)